Two angles are complementary if the sum of their angles is 90 degrees. Suppose we have 2 complementary angle where one angle is ten more than three times the other angle. Determine the measure of the 2 angles.
Answer by checkley71(8403)

x+3x+10=90
4x=90-10
4x=80
x=20 answer for the smaller angle.
3*20+10=60+10=70 for the larger angle.
proof
20+70=90
90=90